Post Go back to editing

AD9371 receive path spikes

Hello,

I have started to see spikes in the data from the AD9371 receiver and would like to understand the cause and eliminate them.

Initially I tested the AD9371 with the default centre frequency of 2.55GHz. The data was captured from JESD into RAM and analysed over of a range of test tone frequencies and amplitudes. The amplitudes were measured in -10dbm intervals down to -100dbm and no-signal. At the lower end the output was noise in the range of +/-10 counts.

Altering the centre frequency to 1.3GHz and re-measuring now shows single sample noise spikes in the input signal when significantly lower than 8-bits (eg/ +/-40 counts). Spikes are in the range -256 to -249 counts, and +248 to +255 appear over about 1% of the samples (example time domain plot of captured test data attached showing problem).

I am surprised that just changing the centre frequency would cause this, however now I am wondering about the LO/PLL and problems with it, and things like decoupling. The hardware is a AD9371 evaluation board.

Any ideas as to the cause of this and how to eliminate the source of the noise?

AD9371-receive-path-spikes.pptx
  • Are you using TES GUI or IIO?

    We did the same with TES GUI and could not reproduce the issue . Can you share the raw I and Q file which you are plotting?

  • Hello,

    I am using TES to generate a profile (header/source C files) to build into a headless (no OS) version of the Mykonos SW (version 1.5.2-3566) which produces my own embedded control SW based on the ADI supplied SW driver. I have my own Xilinx FPGA which hosts the JESD interface.

    I have attached the raw I and Q file, and the excel file which imported this data and has the graph of the I and Q data. Interestingly I noticed all the spikes appear to be on the Q samples...

    IQ-data-2K-samples.zip
  • Does the spike problem resolve if you disable calibration: Rx QEC (quadrature error correction) and/or Rx LOL (LO leakage correction).

  • No it does not resolve. I ran a before (with existing cals) and after test (removing the initial calibrations for RX_LO_DELAY, RX_QEC_INIT). I can see same characteristic spikes on both outputs as before.

  • Is the LO locked at 1.3GHz ? Have you tried any other frequency ?

    Is your TX working at 1.3GHz.or Tx as well has some problem. ? 

  • The LO appears locked at 1.3GHz and yes both Tx and Rx are at the same frequency.

    I have not seen this problem on Tx, just Rx.

    Also previously we were using the default values from the TES tool (2.55GHz an 2.6GHz).

    This setup does not suffer from this spikes problem on Rx.

    So on face value I would say the change prior to the problem occurring was the LO frequency change.

    But why/how would this cause these noise spikes on I-image at low amplitude (<count of 256)?

  • If you use TES GUI SW are you seeing this issue ? (Just to rule out HW related issue).

    If with GUI you are not seeing any issue , problem will be with you NO-OS software . You mentioned above that calibration has no impact to this issue. Disabled or enabled spikes are there.

    Hope you are following the Initialization sequence as mentioned in UG-992. Please verify same.